# Combinatorial Selection with Costly Information

### v1

@@ A.1 Amortized Surrogate Costs for Markov Chains
    Promise of Payment. Whenever one of the Markov chains $\mathcal{M}_{i}$ gets advanced to a state $s_{i}\in S_{i}$ such that $b_{s_{i}^{\prime}t_{i}}>0$ for some terminal state $t_{i}\in T(s_{i})$ and some ancestor state $s_{i}^{\prime}$ of $s_{i}$, the algorithm will immediately advance $s_{i}$.
@@
We note the the promise of payment property states that whenever nature realizes a terminal state that shares a fraction of the cost of some of the actions leading to it, and the algorithm takes one of these actions, then it will ensure that this state is accepted. This ensures that the algorithm will pay back (in expectation) the promised cost shares of all the actions it takes.
Water Filling Amortization.
We proceed to algorithmically define water filling amortization. Fix any Markov chain $\mathcal{M}=(S,\sigma,A,c,\mathcal{D},V,T)$ and for each state $s$, let $\mathrm{lev}(s)$ denote its level; that is, the maximum length of a path from $s$ to some terminal state in $T(s)$. By definition, the level of all terminal states is 00 and the level of the starting state $\sigma$ equals the horizon $H$ of the Markov chain. The algorithmic construction of water filling amortization is given in Algorithm 4.
Input: A Markov chain $\mathcal{M}=(S,\sigma,A,c,\mathcal{D},V,T)$.
1 For each terminal state $t$, initialize $w(t):=v(t)$.
2 for $\ell=1$ to $H$ do
3       for each state $s\in S$ with $\mathrm{lev}(s)=\ell$ do
4             Order $T(s)$ by increasing current costs: $w(t_{1})\leq w(t_{2})\leq\cdots\leq w(t_{m})$.
5             Set $b_{st}:=0$ for all $t\in T(s)$.
6             Initialize $i:=0$ and let $w(t_{m+1}):=\infty$.
7             while $c(s)\cdot p(s)>\sum_{t\in T(s)}p(t)\cdot b_{st}$ do
8                   Increment $i:=i+1$.
9                   Compute threshold $T:=\big{(}c(s)p(s)-\sum_{t\in T(s)}p(t)b_{st}\big{)}/\big{(}\sum_{j\leq i}p(t_{j})\big{)}$.
10                   Set $b_{st_{i}}:=\min\{T\;,\;w(t_{i+1})-w(t_{i})\}$ and $w(t_{i}):=w(t_{i})+b_{st_{i}}$.
11                   For all $j<i$, update $b_{st_{j}}:=b_{st_{j}}+b_{st_{i}}$ and $w(t_{j}):=w(t_{j})+b_{st_{i}}$.
12
13             end while
14
15       end for
16
17 end for
Output: The water filling amortization $\{b_{st}\}$ for all $s\in S$ and $t\in T(s)$.
Algorithm 4 Water Filling Amortization
The algorithm visits the states of the Markov chain in increasing level order. When considering state $s$, it needs to determine cost shares $b_{st}\geq 0$ such that $c(s)p(s)=\sum_{t\in T(s)}p(t)b_{st}$. It starts by distributing its cost to the terminal state $t_{1}$ of minimum current cost $w(t_{1})$ until it matches the current cost of the terminal state $t_{2}$ that has the second minimum cost; after that point, the two terminals get grouped together and are increased simultaneously from now on. Computing the threshold $T$ at every iteration ensures that we will never allocate more cost than necessary.
We note that an equivalent way to define the amortization at state $s$ is to compute the index $g_{s}\geq 0$ that satisfies
$c(s)p(s)=\sum_{t\in T(s)}p(t)\cdot(g_{s}-w(t))^{+}$
and then set $b_{st}:=(g_{s}-w(t))^{+}$ for each $t\in T(s)$. A pictorial representation of water filling amortization is shown in Figure 8.
Figure 8: For a state $s$ with $T(s)=\{t_{1},t_{2},t_{2}\}$, the water filling algorithm will find the (unique) $g_{s}\geq 0$ for which the highlighted area becomes $c(s)$. Then, it sets $b_{st}=(g_{s}-w(t))^{+}$ and updates $w(t)=\max(g_{s},w(t))$ for all $t\in T(s)$.
Optimality of Water Filling Index Policy.
Finally, we will combine everything to argue about the optimality of the water filling index policy, establishing Theorem 3.2. See 3.2
@@
From Corollary A.1, it suffices to argue that the water filling index policy, paired with the water filling amortization, satisfies both surrogate optimality and promise of payment.
We begin with surrogate optimality. By definition, the water filling index policy advances the chain of minimum water filling index; recall that the index of a state corresponds to the minimum surrogate cost among its terminal states. This ensures that while there is potential for some chain $\mathcal{M}_{i}$ to realize the terminal state of minimum surrogate cost, the algorithm will keep advancing it. In other words, this algorithms ends up greedily accepting Markov chains with respect to the surrogate cost of their realized terminal states. Paired with the fact that a minimum cost basis of a matroid is always obtained by greedily adding the cheaper feasible element, this establishes surrogate optimality of the water filling index policy.
Next, we show that promise of payment also holds. Say that at any point the algorithm advances Markov chain $\mathcal{M}_{i}$ with index $I_{i}$ and reaches a state $s_{i}\in S_{i}$ such that $b_{s_{i}^{\prime}t_{i}}>0$ for some terminal state $t_{i}\in T(s_{i})$ and some ancestor $s_{i}^{\prime}$ of $s_{i}$. Since $\mathcal{M}_{i}$ was advanced, $I_{i}$ was the minimum index among all Markov chains; if we can prove that the index at state $s_{i}$ is also $I_{i}$, then the algorithm will keep advancing $\mathcal{M}_{i}$ and the promise of payment property will hold. Since $b_{s_{i}^{\prime}t_{i}}>0$ for some terminal state $t_{i}\in T(s_{i})$ and some ancestor $s_{i}^{\prime}$ of $s_{i}$, we know that when $s^{\prime}_{i}$ was being amortized, $t_{i}$ was the terminal state in $T(s^{\prime}_{i})$ with minimum current surrogate cost. By definition of the water filling amortization, this implies that $t_{i}$ will continue to have the minimum surrogate cost among all the terminal states in $T(s^{\prime}_{i})$ for the rest of the amortization process, implying that the index of all states in the path from $t_{i}$ to $s^{\prime}_{i}$ (including state $s_{i}$) will be the same. ∎
A.2 Amortized Surrogate Costs for MDPs
@@
encode all the randomness in the algorithm’s run except from the realizations of $\mathcal{M}_{i}$. The key observation is that conditioned on $\mathcal{R}^{-i}$, all the actions that the algorithm takes in MDP $\mathcal{M}_{i}$ are predetermined; in other words, the algorithm’s trajectory on $\mathcal{M}_{i}$ is fully described by some committing policy $\pi_{i}=\pi_{i}(\mathcal{R}^{-i})\in\mathcal{C}\left(\mathcal{M}_{i}\right)$. Thus, we have that
@@
where $cost(\pi_{i})$ is the expected cost of the algorithm, following the committing policy $\pi_{i}=\pi_{i}(\mathcal{R}^{-i})$, on MDP $\mathcal{M}_{i}$. The proof is the completed by the following generalization of Lemma 3.1, which is enabled from Lemma 3.3.
@@
Fix any MDP $\mathcal{M}$ and any committing policy $\pi\in\mathcal{C}\left(\mathcal{M}\right)$. Then, the expected cost of any algorithm following $\pi$ on $\mathcal{M}$ will be at least
@@
Notice that the above claim doesn’t depend on the underlying CICS instance $\mathbb{I}$; it simply states that conditioned on running a committing policy on some MDP and accepting a terminal state, the expected total cost spent on this MDP is lower bounded by the surrogate cost. Since this lower bound applies to all committing policies $\pi\in\mathcal{C}\left(\mathcal{M}\right)$, coupled with equation (12), it directly implies inequality (11), as
@@
Proof of 3.
Notice that since we are committing to running policy $\pi\in\mathcal{C}\left(\mathcal{M}\right)$ on $\mathcal{M}$, we are essentially running some algorithm on the Markov chain $\mathcal{M}^{\pi}$. From Lemma 3.1 and Theorem 3.2, we know that the contribution of Markov chain $\mathcal{M}^{\pi}$ to the total cost will be at least
@@
Let $\mathcal{M}=(S,\sigma,A,c,\mathcal{D},V,T)$ and recall that we use $S_{\pi}$ and $T_{\pi}$ to denote the state space and terminal states of $\mathcal{M}^{\pi}$ and $p_{\pi}$ to denote the implied distribution over $T_{\pi}$. Finally, we use $c(s)$ to denote the cost of the unique action that $\pi$ chooses at state $s\in S_{\pi}$. By definition of the water filling surrogate cost of a Markov chain, we have
$\operatorname{E}\left[W^{*}_{\mathcal{M}^{\pi}}\right]$ $=\sum_{t\in T_{\pi}}p_{\pi}(t)\cdot W^{*}_{\mathcal{M}^{\pi}}(t)=\sum_{t\in T_{\pi}}p_{\pi}(t)\cdot\bigg{(}v(t)+\sum_{s\in S_{\pi}:t\in T_{\pi}(s)}b^{*}_{st}\bigg{)}$
where $b^{*}_{st}$ are non-negative cost shares satisfying $\sum_{t\in T_{\pi}(s)}p_{\pi}(t)b^{*}_{st}=p_{\pi}(s)c(s)$ for all $s\in S_{\pi}$.
From Lemma 3.3, there exists an amortized cost function $\rho(\cdot)$ over the terminal states $T$ of $\mathcal{M}$ and a set of non-negative cost shares $\{b\}_{st}$ that satisfy action independence, cost sharing and cost dominance. Using these, we have
$\operatorname{E}\left[W^{*}_{\mathcal{M}^{\pi}}\right]$ $=\sum_{t\in T_{\pi}}p_{\pi}(t)\cdot\bigg{(}v(t)+\sum_{s\in S_{\pi}:t\in T_{\pi}(s)}b^{*}_{st}\bigg{)}$
$=\sum_{t\in T_{\pi}}p_{\pi}(t)\cdot\bigg{(}\operatorname{E}\left[\rho(t)\right]-\sum_{s\in S_{\pi}:t\in T_{\pi}(s)}b_{st}+\sum_{s\in S_{\pi}:t\in T_{\pi}(s)}b^{*}_{st}\bigg{)}$ (Cost Sharing)
$=\sum_{t\in T_{\pi}}p_{\pi}(t)\cdot\operatorname{E}\left[\rho(t)\right]+\sum_{s\in S_{\pi}}\sum_{t\in T_{\pi}(s)}p_{\pi}(t)\cdot(b^{*}_{st}-b_{st})$
$=\sum_{t\in T_{\pi}}p_{\pi}(t)\cdot\operatorname{E}\left[\rho(t)\right]+\sum_{s\in S_{\pi}}\bigg{(}p_{\pi}(s)c(s)-\sum_{t\in T_{\pi}(s)}p_{\pi}(t)\cdot b_{st}\bigg{)}$
$\geq\sum_{t\in T_{\pi}}p_{\pi}(t)\cdot\operatorname{E}\left[\rho(t)\right]$ (Cost Dominance)
$=\operatorname{E}\left[W^{*}_{\mathcal{M}}\right].$ (Action Independence)
A.3 Second Order Stochastic Dominance
Finally, in this section we provide a proof for Lemma 3.5. Once again, we note that this result is standard; here, we provide our own constructive proof for the sake of completeness and building intuition on how the water filling surrogate cost of an MDP is obtained. For simplicity, we refer to Lemma 3.5 as the Stochastic Dominance Lemma, henceforth SDL. See 3.5
@@ A.3 Second Order Stochastic Dominance
Appendix B Omitted Proofs from Section 4
In this section, we present the omitted proof of Theorem 4.1 from Section 4. We restate all our results for the reader’s convenience. See 4.1
@@ Appendix B Omitted Proofs from Section 4
where $\tilde{W}_{1}=\alpha\cdot W^{*}_{\mathcal{M}_{1}}$ and $\tilde{W}_{i}:=W^{*}_{\mathcal{M}_{i}^{\pi_{i}}}$ for all $i\neq 1$. Thus, we have substituted $W^{*}_{\mathcal{M}_{1}^{\pi_{1}}}$ with $\alpha\cdot W^{*}_{\mathcal{M}_{1}}$. The proof is completed by repeating the same process for all other $i\neq 1$. ∎
Appendix C Omitted Proofs from Section 5
In this chapter of the appendix, we present all omitted proofs from Section 5, in particular the proof of Lemma 5.2. We restate all our results for the reader’s convenience.

### v2

@@ D.1 Amortized Surrogate Costs for Markov Chains
    Promise of Payment. Whenever one of the Markov chains $\mathcal{M}_{i}$ gets advanced from a state $s_{i}\in S_{i}$ such that $b_{s_{i}\tau_{i}}>0$ for some $\tau_{i}\in\Upsilon(s_{i})$, the algorithm will keep advancing $\mathcal{M}_{i}$ as long as it’s trajectory evolves according to $\tau_{i}$.
@@
Promise of payment states that whenever $b_{s\tau}>0$ and the algorithm advances $s$, it will ensure that while it is possible for $\tau$ to be realized, the corresponding Markov chain will continue to be advanced; in other words, that $\operatorname{Pr}\left[I(s)\right]\cdot\operatorname{Pr}\left[R(\tau)|R(s)\right]=\operatorname{Pr}\left[A\cap R(\tau)\right]$. We note that achieving any of these properties individually is trivial; fully advancing all Markov chains and picking the feasible set of trajectories of minimum total surrogate cost achieves surrogate optimality, and either advancing a Markov chain fully or not advancing it at all achieves promise of payment. The water filling amortization is specifically defined so that the corresponding index policy achieves both of these properties, simultaneously: See 3.2
@@
It suffices to argue that the water filling index policy, paired with the water filling amortization, satisfies both surrogate optimality and promise of payment. We begin with surrogate optimality. By definition, the water filling index policy advances the chain of minimum water filling index; recall that the index of a state corresponds to the minimum surrogate cost among all trajectories passing through it. This ensures that while there is potential for some chain $\mathcal{M}_{i}$ to realize the trajectory of minimum surrogate cost, the algorithm will keep advancing it. In other words, this algorithms ends up greedily accepting Markov chains with respect to the surrogate cost of their realized trajectories. Paired with the fact that a minimum cost basis of a matroid is always obtained by greedily adding the cheaper feasible element, this establishes surrogate optimality of the water filling index policy.
Next, we show that promise of payment also holds. Say that at any point the algorithm advances Markov chain $\mathcal{M}_{i}$ at state $s_{i}$ with index $I_{i}(s_{i})$ and there exists some trajectory $\tau_{i}\in\Upsilon(s_{i})$ such that $b_{s_{i}\tau_{i}}>0$. We need to argue that as long as the state of $\mathcal{M}_{i}$ evolves according to $\tau_{i}$, the algorithm will keep advancing it. Since $\mathcal{M}_{i}$ was advanced by the water filling index policy, this implies that $I_{i}(s_{i})$ was the minimum index among the current states of all Markov chains; thus, a sufficient condition is to show that $I_{i}(s^{\prime}_{i})\leq I_{i}(s_{i})$ for all $s^{\prime}_{i}\in\tau_{i}$, as then all the states in the trajectory will continue to have the minimum index and thus will be advanced if realized. Since $b_{s_{i}\tau_{i}}>0$, we know by definition of the water filling amortization that immediately after $s_{i}$ got amortized, $\tau_{i}$ had the minimum surrogate cost across all trajectories in $\Upsilon(s_{i})$. As states are considered in a bottom-up order, this implies that any subsequent steps of the amortization will maintain that the trajectory passing through $s_{i}$ that has the minimum surrogate cost ends with $\tau_{i}$. This immediately gives us the proof, by definition of the indices and $s^{\prime}_{i}\in\tau_{i}$. ∎
D.2 Amortized Surrogate Costs for MDPs
@@
encode all the randomness in the algorithm’s run except from the realizations of $\mathcal{M}_{i}$. The key observation is that conditioned on $\mathcal{R}^{-i}$, all the actions that the algorithm takes in MDP $\mathcal{M}_{i}$ are predetermined; in other words, the algorithm’s trajectory on $\mathcal{M}_{i}$ is fully described by some deterministic commitment denoted $\pi_{i}=\pi_{i}(\mathcal{R}^{-i})\in\mathcal{C}\left(\mathcal{M}_{i}\right)$. Thus, we have that
@@
where $cost(\pi_{i})$ is the expected cost of the algorithm, following the commitment $\pi_{i}=\pi_{i}(\mathcal{R}^{-i})$, on MDP $\mathcal{M}_{i}$. The proof is the completed by the following generalization of Lemma 3.1, which is enabled from Lemma 3.5.
@@
Fix any MDP $\mathcal{M}$ and any deterministic commitment $\pi\in\mathcal{C}\left(\mathcal{M}\right)$. Then, the expected cost of any algorithm following $\pi$ on $\mathcal{M}$ will be at least
@@
Notice that the above claim doesn’t depend on the underlying CICS instance $\mathbb{I}$; it simply states that conditioned on running a committing policy on some MDP and accepting a terminal state, the expected total cost spent on this MDP is lower bounded by the surrogate cost. Since this lower bound applies to all commitments $\pi\in\mathcal{C}\left(\mathcal{M}\right)$, coupled with equation (9), it directly implies inequality (8), as
@@
Proof of 4
. Notice that since we are committing to running policy $\pi\in\mathcal{C}\left(\mathcal{M}\right)$ on $\mathcal{M}$, we are essentially running some algorithm on the Markov chain $\mathcal{M}^{\pi}$. From Lemma 3.1 and Theorem 3.2, we know that the contribution of Markov chain $\mathcal{M}^{\pi}$ to the total cost will be at least
@@
Let $\mathcal{M}=(S,\sigma,A,c,\mathcal{D},V,T)$ and recall that we use $S_{\pi}$ and $\Upsilon_{\pi}$ to denote the state space and trajectory set of $\mathcal{M}^{\pi}$ and $p_{\pi}$ to denote the implied distribution over $\Upsilon_{\pi}$. Finally, we use $c(s)$ to denote the cost of the unique action that the deterministic commitment $\pi$ chooses at state $s\in S_{\pi}$. By definition of the water filling surrogate cost of a Markov chain, we have
$\operatorname{E}\left[W^{*}_{\mathcal{M}^{\pi}}\right]$ $=\sum_{\tau\in\Upsilon_{\pi}}p_{\pi}(\tau)\cdot\rho^{*}_{\mathcal{M}^{\pi}}(\tau)=\sum_{\tau\in\Upsilon_{\pi}}p_{\pi}(\tau)\cdot\bigg{(}v(\tau)+\sum_{s\in\tau}b^{*}_{s\tau_{s}}\bigg{)}$
where $b^{*}_{s\tau}$ are non-negative cost shares satisfying $\sum_{\tau\in\Upsilon_{\pi}(s)}p_{\pi}(\tau)b^{*}_{s\tau}=p_{\pi}(s)c(s)$ for all $s\in S_{\pi}$. Since $\pi$ is deterministic, we know from Lemma 3.5 that there exists an amortized cost function $\rho^{\pi}(\cdot)$ over the trajectories in $\Upsilon_{\pi}$ and a set of cost shares $\{b^{\pi}_{s\tau}\}$ that will satisfy action independence, cost sharing and cost dominance. Using these, we have
$\operatorname{E}\left[W^{*}_{\mathcal{M}^{\pi}}\right]$ $=\sum_{\tau\in\Upsilon_{\pi}}p_{\pi}(\tau)\cdot\bigg{(}v(\tau)+\sum_{s\in\tau}b^{*}_{s\tau_{s}}\bigg{)}$
$=\sum_{\tau\in\Upsilon_{\pi}}p_{\pi}(\tau)\cdot\bigg{(}\operatorname{E}\left[\rho^{\pi}(\tau)\right]-\sum_{s\in\tau}b^{\pi}_{s\tau_{s}}+\sum_{s\in\tau}b^{*}_{s\tau_{s}}\bigg{)}$ (Cost Sharing)
$=\sum_{\tau\in\Upsilon_{\pi}}p_{\pi}(\tau)\cdot\operatorname{E}\left[\rho^{\pi}(\tau)\right]+\sum_{\tau\in\Upsilon_{\pi}}\sum_{s\in\tau}p_{\pi}(\tau)(b^{*}_{s\tau_{s}}-b^{\pi}_{s\tau_{s}})$
$=\operatorname{E}\left[W^{*}_{\mathcal{M}}\right]+\sum_{s\in S_{\pi}}\sum_{\tau\in\Upsilon_{\pi}(s)}p_{\pi}(\tau)(b^{*}_{s\tau}-b^{\pi}_{s\tau})$ (Action Independence)
$=\operatorname{E}\left[W^{*}_{\mathcal{M}}\right]+\sum_{s\in S_{\pi}}p_{\pi}(s)c(s)-\sum_{\tau\in\Upsilon_{\pi}(s)}p_{\pi}(\tau)b^{\pi}_{s\tau}$
$\geq\operatorname{E}\left[W^{*}_{\mathcal{M}}\right].$ (Cost Dominance)
D.3 Second Order Stochastic Dominance
Finally, in this section we provide a proof for Lemma 3.6. Once again, we note that this result is standard; here, we provide our own constructive proof for the sake of completeness and building intuition on how the water filling surrogate cost of an MDP is obtained. For simplicity, we refer to Lemma 3.6 as the Stochastic Dominance Lemma, henceforth SDL. See 3.6
@@ D.3 Second Order Stochastic Dominance
Appendix E Omitted Proofs from Section 4 (Local Approximation)
In this section, we present the omitted proof of Theorem 4.1 from Section 4. See 4.1
@@ Appendix E Omitted Proofs from Section 4 (Local Approximation)
where $\tilde{W}_{1}=\alpha\cdot W^{*}_{\mathcal{M}_{1}}$ and $\tilde{W}_{i}:=W^{*}_{\mathcal{M}_{i}^{\pi_{i}}}$ for all $i\neq 1$. Thus, we have substituted $W^{*}_{\mathcal{M}_{1}^{\pi_{1}}}$ with $\alpha\cdot W^{*}_{\mathcal{M}_{1}}$. By repeating the same process for $i=2,3,\cdots,n$, we obtain that
$\operatorname{E}\left[\min_{S\in\mathcal{F}}\sum_{i\in S}W^{*}_{\mathcal{M}_{i}^{\pi_{i}}}\right]\leq\operatorname{E}\left[\min_{S\in\mathcal{F}}\sum_{i\in S}\alpha\cdot W^{*}_{\mathcal{M}_{i}}\right]$
and thus
$\frac{\operatorname{E}\left[\min_{S\in\mathcal{F}}\sum_{i\in S}W^{*}_{\mathcal{M}_{i}^{\pi_{i}}}\right]}{\operatorname{E}\left[\min_{S\in\mathcal{F}}\sum_{i\in S}W^{*}_{\mathcal{M}_{i}}\right]}\leq\alpha.$
From Theorem 3.2 the nominator is precisely the cost of the optimal committing policy for $\mathbb{I}$ under commitments $\mathcal{P}=(\pi_{1},\cdots,\pi_{n})$ and from Theorem 3.4, the denominator is a lower bound on $\mathrm{OPT}(\mathbb{I})$. Thus, we have $\mathrm{CG}(\mathbb{I})\leq\alpha$ as desired.
Appendix F Omitted Proofs from Section 5 (Partial Inspection PB)
F.1 Proof of Lemma 5.2